OpenStudy (anonymous):

If h(x) = -3x - 10, find h(-3).

OpenStudy (anonymous):

Again, plug in -3 for x. You get h(-3) = -3(-3)-10 = 9-10 = -1

jhonyy9 (jhonyy9):

h(-3)=-1